# Synthetic Biology Meets Neuromorphic Computing: Towards a Bio-Inspired Olfactory Perception System

### v1

@@ Abstract
In this study, we explore how the combination of synthetic biology, neuroscience modeling, and neuromorphic electronic systems offers a new approach to creating an artificial system that mimics the natural sense of smell. We argue that a co-design approach offers significant advantages in replicating the complex dynamics of odor sensing and processing. We investigate a hybrid system of synthetic sensory neurons that provides three key features: a) receptor-gated ion channels, b) interface between synthetic biology and semiconductors and c) event-based encoding and computing based on spiking networks. This research seeks to develop a platform for ultra-sensitive, specific, and energy-efficient odor detection, with potential implications for environmental monitoring, medical diagnostics, and security.
@@ 1 Introduction
Biological olfaction is a remarkably efficient sensory system, surpassing engineered chemical detection methods through redundant neural circuits that ensure robust responses [1], combinatorial coding that enables the detection of thousands of odors [2, 3], and exceptional sensitivity  [4]. These advantages have inspired extensive efforts to replicate biological olfaction, leading to significant improvements in artificial olfaction. Today’s electronic noses (e-noses) and artificial olfactory biosensors employ arrays of chemosensors and advanced pattern recognition algorithms, enabling odor classification in an impressive variety of fields, from biomedical signal analysis [5] and environmental monitoring [6], to food safety [7], agricultural productivity [8], and security [9].
Despite these notable achievements, e-nose technologies continue to face several key challenges that have slowed their broader adoption. These include: (i) sensor drift, which deteriorates performance over time [10]; (ii) cross-sensitivity, which leads to reduced specificity when sensors respond to multiple volatile compounds [11, 12]; (iii) significant power consumption [13], hampering the feasibility of portable or embedded deployments [12]; (iv) limited computational capacity for interpreting the complex combinatorial patterns inherent in real-world odors [14]; and (v) insufficient adaptability, as current systems cannot dynamically refine their responses from experience, unlike their biological counterparts [4]. Observing how biological olfaction circumvents these problems suggests a promising model for the next generation of artificial olfactory devices, particularly given its robust signal processing capabilities, cross-selective receptor integration, and the ability to learn from repeated exposure [15, 14, 16].
@@
Crucially, olfactory receptors do not respond to each odorant in isolation; instead, volatile compounds compete for overlapped receptor sites through a process often described as odor occlusion [19, 20]. These competitive interactions can be highly non-linear: one odorant can amplify receptor activity, while another can dampen it, leading to outcomes that defy simple additive models [20]. Such receptor-level dynamics introduce yet another layer of complexity, making biological olfaction remarkably adept at navigating real-world odor spaces. Unlocking these principles in engineered systems promises to push artificial olfaction beyond its current limits, paving the way for more robust, adaptable, and context-aware sensing technologies.
@@ 1.1 Background & state-of-the-art
Metal oxide sensors, while effective for chemical detection, face significant drift that compromises their long-term reliability and applicability in the real world [27, 28]. This drift manifests as temporal shifts in sensor response under identical working conditions, primarily caused by chemical and physical interactions at the sensing film microstructure, as well as external factors such as temperature and humidity variations. For example, recent analysis of the widely used gas sensor dataset by Vergara et al. [29] revealed how sensor drift can significantly affect recordings, hindering realistic assessment of sensor performance [28]. Although traditional approaches attempt to address drift through regression models and environmental compensation [30, 31], bio-inspired neuromorphic implementations offer promising alternatives through spike-based encoding [32] and adaptive learning mechanisms that could potentially provide stability and resistance to drift [33]. These neuromorphic approaches aim to reduce the complexity of the data and implement adaptive algorithms that can follow changing conditions, potentially offering a more robust solution to the persistent challenge of sensor drift.
@@
Building on these concepts, integrated analog on-chip learning circuitry has demonstrated efficacy in odor detection and classification. For example, Covington et al. in [24] used an 80 element chemoresistive microsensor array combined with a microfluidic package, while Tang et al. in [26] relied on commercially available sensors and a traditional signal processing pipeline in a microprocessor to distinguish three fruit fragrances. In a related work [25], Bernabei et al. utilized a large array of chemosensors with slightly varied sensor elements to emulate the overlap selectivity characteristic of biological olfaction, generating rich combinatorial responses. More recently, Rastogi et al. [34] introduced a neuromorphic analog e-nose that directly encodes the gas concentration using metal oxide (MOx) sensors and an analog front-end that transforms sensor responses into spike-based representations. Their approach leverages the time difference between spikes in two distinct processing pathways to infer gas concentration, reducing data redundancy, and improving power efficiency.
@@
However, current polymer and metal oxide-based sensing elements remain relatively simple compared to biological olfactory systems, facing challenges with odor specificity and stability [11, 35]. Although these artificial sensors can detect certain volatile compounds, they often require high operating temperatures and exhibit limited effectiveness in discriminating complex molecular mixtures [36]. Moreover, their continuous sampling needs and high power consumption stand in stark contrast to the low-power, event-driven nature of biological olfaction [4]. These fundamental discrepancies in functionality and energy efficiency highlight the limitations of conventional chemosensors in achieving the richness and adaptability seen in biological odor detection.
@@
Notably, some more recent works have demonstrated a more complete reconstitution of the olfactory system, in particular, into bio-mimetic environments, for example membrane bound vesicles formed by membrane extraction, or virus-like particles [39, 40, 41, 42]. These approaches leverage the fact that insect olfaction is largely based on multivalent interactions between the conserved co-receptor Orco and the odorant binding subunits (odorant receptors, ORs), which together form a ligand-gated ion channel. Khadka et al. have demonstrated successful insertion of multiple odorant receptors into surface-anchored lipid vesicles. Yamada et al. [42] have demonstrated parts per billion sensitivity of an Orco-OR complex inserted into a lipid bilayer to octenol, a biomarker in human breath. Detection of femtomolar concentrations of VUAA1, E2-hexenal, 4-ethylguaiacol molecules was demonstrated using odorant receptors Or10a, Or22a, and Or71a [41].
@@
The signal restoration challenge is being recognized as one of the central problems in neuromorphic olfaction [43, 44, 4]. In fact, for more than 20 years, researchers have concluded that the speed and accuracy of odor recognition in insects cannot be attributed solely to the performance of olfactory receptors due to their slow time constant and high variability [45]. The requirement to correctly classify odor concentrations poses additional demand on the analyzing algorithms, as do signal occlusion and noise effects. Thus, several models of signal processing downstream of the sensor have been investigated, e.g. [46, 47, 48, 49, 16, 45, 50, 51, 52, 53]. Methods for odor signal analysis can be broadly categories into two classes: data analysis tools (e.g., principal component analysis, support vector machines) and biorealistic network models. While data analysis tools have been used for odor classification since many years [46, 47], biologically realistic network models are now studied with the goal of reaching the accuracy and efficiency of the animal olfactory system [50, 51, 16, 52]. Many of these models perform classification using a spatiotemporal code while learning in an unsupervised fashion. Additionally, some studies attribute the efficiency of natural systems to two main intrinsic mechanisms of biorealistic neural networks (i.e., Spiking Neural Network (SNN)): i) inhibition (which triggers a winner-takes-all competition [48, 50, 16]) and ii) temporal integration (due to neuron and network dynamics) [45, 16]. Recent work has also pursued the implementation of learning from teaching signals (supervised and reinforcement learning) in models of the insect olfactory system, in particular for associative learning [53, 54].
@@
Several hardware implementations have been demonstrated in digital Field Programmable Gate Arrays (FPGA) hardware [50], digital neuromorphic devices [16] and analog neuromorphic chips [23, 48, 55, 34]. Recently, a bio-realistic network [16], based on the architecture of the mammalian olfactory bulb, was implemented on the Loihi neuromorphic digital chip, demonstrating the benefits of employing an algorithm based on online learning and spike timing. However, the system was trained using a separately recorded data set of pure chemicals [29], therefore lacking integration with real chemosensors, and criticism has been raised regarding the generalizability of the network model [28].
The neuromorphic olfactory sensor literature indicates that there is considerable scope for improving neuromorphic chemosensory systems. There are several limitations in many of the current implementations, such as the lack of signal conditioning, the lack of integration of sensing and decoding in a single platform, high power consumption, and the poor sensitivity of the artificial sensor array. So far, the only truly integrated neuromorphic sensing platform for olfaction has been developed by Koickal et al. in 2006 [56, 23]. To our knowledge, this platform did not leave the proof-of-concept stage, and subsequent work by the authors has used conventional processing hardware [50]. This work is focused on integrating hardware realizations of all stages into one singular system.
@@ 2 Co-design of synthetic biology, bio-inspired models, and neuromorphic systems
A comprehensive approach is crucial in developing advanced neuromorphic chemosensory systems that effectively integrate sensor arrays, decoding algorithms, and their physical realization [57]. In this work, we present a novel methodology and preliminary results on how to move beyond the State-Of-The-Art (SOTA) by addressing the lack of codesign between these three components. Unlike previous chemical detection systems inspired by the olfactory pathway, which have treated sensing [58, 47], signal processing [59, 60], and neuromorphic computation [16] as separate components, our approach integrates these aspects to tackle the challenge of chemosensory perception in a comprehensive way. In this work, we introduce a new concept to couple the depolarization of synthetic cells made from biological materials directly with CMOS electronic circuits designed for spike generation, we call these new class of bio-electronic devices hybrid synthetic sensory neurons, and we depict them in Figure 1.
@@
In doing so, our framework harnesses the advantages of co-design in a joint environment. Here, the three disciplines (neuroscience, synthetic biology, neuromorphic engineering) are developed together and connected to each other, as depicted in Figure 1:
@@
Our approach is not without challenges, as individual parts need to be optimized and tuned together but must still allow for the necessary flexibility. This flexibility is needed as knowledge and research about the individual parts progresses simultaneously. The parallel approach is, on one hand, due to constraints of scientific funding—collaborative grants typically work on a common timeline from proposal to project end. However, parallelized development is also, in general, desired for time efficiency. As such, the collaborative tools we develop, particularly those that bridge data-driven synthetic biology with electrical and information engineering, not only represent standalone contributions to each field, but might also be useful in other applications.
3 Results & Discussions
@@ 3.1 Insect odor receptors provide a modular sensing platform
Insect olfaction is largely based on membrane-bound protein complexes consisting of an odorant receptor (OR) and a co-receptor, Orco. While ORs diverge in their sequence and odor specificity between the multitude of insects and the ecological niches they inhabit, the co-receptor Orco is highly conserved across insect species. Several experiments show that it is possible to create functional receptor complexes by combining Orco of one species with a foreign OR [61, 62]. From an engineering point of view, this suggests the possibility of a modular platform in which the base assembly of an Orco-based sensing element allows the swapping of ORs with different response profiles to target various compounds. To demonstrate this modularity, we set out to computationally assemble Orco-OR complexes using AlphaFold modeling [63, 64].
@@
Figure 2: AlphaFold predictions suggest hundreds of possible stable complexes between D. melanogaster Orco and foreign odor receptors (OR). a1: AlphaFold ipTM score for Orco-OR complexes formed from D. melanogaster Orco combined with OR of different insect species (black), and combined with D. melanogaster (Dm)  OR (positive control, red) and M. musculus OR (negative control, grey). a2: Interface solvation energy prediction for Dm with foreign OR above the ipTM cut-off of 0.6. Dashed grey lines indicates cut-off thresholds. b1: Dm Orco (red) dimerization with M. sexta OR4 resulted in an ipTM score above 0.6 (dark grey) while the dimerization with L. migratoria OR113 led to a score below 0.6 (grey). b2:  Dm Orco and B. mori OR15 complex with favorable solvation energy ($>-5$ kcal/mol), whereas the dimerization with C. quinquefasciatus OR151 resulted in a low ($<-5$ kcal/mol) solvation energy. Varying number of bonds can be observed in the anchor domain (brackets). Green dashed lines indicate electrostatic interactions, while dashed blue lines represents hydrogen bonds.
@@
Specifically, we selected the Orco variant from the D. melanogaster (Dm) fruit fly and considered a total of 348 Orco-OR complexes, including non-Dm ORs. Each predicted Orco-OR structure was evaluated using the interface-predicted template modeling (ipTM) score, with higher values indicating greater confidence in the prediction accuracy of the protein-protein interface. An empirical cut-off point of 0.6 was chosen (Figure 2a1) [65, 64]. As expected, the positive control, consisting only of Orco-OR complexes from the same species (Dm), yielded mostly acceptable modeling outputs, as reflected in ipTM scores above 0.6. In contrast, the negative control, which used the mammalian OR from M. musculus, a receptor not expected to form a complex with Orco, did not produce any acceptable models (Figure 2a1). Interestingly, for Dm Orco and ORs from foreign species, we found a large fraction of potentially accurate models with ipTM scores above the 0.6 cut-off. This suggests that AlphaFold models are relevant for understanding the structures of Dm Orco in complex with ORs from foreign species.
@@
To further evaluate the structures, we considered a second metric: interface solvation energy, which provides insight into the stability of the non-covalent binding between Orco and OR (Figure 2a2). More negative values indicate a more favorable binding interface [66]. We found that the obtained values for foreign OR were similar to those of the native Dm Orco-OR complexes (Figure 2b2), indicating comparable complex stability even for foreign ORs. Furthermore, we manually inspected the most favorable predictions for each group. In both cases, the interactions between the two subunits were confined to a segment of helical bundles stabilized by hydrogen bonds and electrostatic interactions located away the membrane interface (Figure 2a2), consistent with the anchor domain and interactions identified in experimental studies [67, 68, 69]. These results suggest that the OR structures generated by AlphaFold are promising candidates for computational exploration of the wide diversity of ORs.
@@
Figure 3: Comparison of ligand binding energies of Dm - Or69a, Or98a, Or19a and Or92a. Each datapoint corresponds to one odor molecule. Left: Terpene subset. Right: All odors from the DoOR dataset with positive neuron response. Dashed line shows linear regression with $R^{2}=0.2$
@@
To demonstrate the potential to build a digital representation of OR-odor interactions, we challenged the generated AlphaFold Orco-OR structures with in-vivo recordings of neuronal activity. We used the DoOR 2.0.1 dataset [70, 71], a comprehensive database of olfactory responses from various measurement techniques such as single sensillum recordings, calcium imaging, and heterologous expression [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 49, 89, 90, 91, 92]. If the predicted OR structures are accurate, we expected that neuronal activity would correlate with the binding of the response-inducing molecule to the OR model. We initially focused on terpenes, a class of carbohydrates with a large natural abundance, because we speculated that these molecules were a feasible target for a computationally relatively simple molecular docking approach. We calculated the binding energy of terpenes, to the generated OR structures using Vina Dock (see Materials and Methods) for a subset of Dm ORs and plotted these against the measured neuronal response from the DoOR dataset (Figure 3, left). We found that the difference in binding energy between responding (> 0) and non-responding neurons ($\leq$ 0) was highly significant (Student’s t-test, $p<3\cdot 10^{-7}$), suggesting that the predicted structures have realistic binding pockets for terpenes. Motivated by this, we investigated the same Dm OR subset using all molecules present in the DoOR dataset. Here, we obtained a weak correlation between neuron activation and binding energy (Figure 3, right). At this point it is unclear whether the scatter can be explained by the uncertainty of the molecular docking method or point to more complex molecular recognition mechanisms. Currently, we are investigating an even larger number of ORs and binding molecules with refined computational approaches. If the responses of olfactory neurons to odors are indeed explained simply by the binding energy of an odor to the OR, the molecular recognition process of insect olfaction can be condensed into a measure of this binding energy. Taken together, these results show the potential to use computational tools to predict OR to odor response.

### v2

@@ Abstract
In this study, we explore how the combination of synthetic biology, neuroscience modeling, and neuromorphic electronic systems offers a new approach to creating an artificial system that mimics the natural sense of smell. We argue that a co-design approach offers significant advantages in replicating the complex dynamics of odor sensing and processing. We propose a hybrid system of synthetic sensory neurons that provides three key features: a) receptor-gated ion channels, b) interface between synthetic biology and semiconductors and c) event-based encoding and computing based on spiking networks. Our approach is validated using simulation-based modelling of the complete sensing and processing pipeline. This research seeks to develop a platform for ultra-sensitive, specific, and energy-efficient odor detection, with potential implications for environmental monitoring, medical diagnostics, and security.
@@ 1 Introduction
Biological olfaction is a remarkably efficient sensory system, surpassing engineered chemical detection methods through redundant neural circuits that ensure robust responses [1], combinatorial coding that enables the detection of multitude of odors [2, 3, 4], and exceptional sensitivity [5]. These advantages have inspired extensive efforts to replicate biological olfaction, leading to significant improvements in artificial olfaction  [6, 7]. Today’s electronic noses (e-noses) and artificial olfactory biosensors employ arrays of chemosensors and advanced pattern recognition algorithms, enabling odor classification in an impressive variety of fields, from biomedical signal analysis [8] and environmental monitoring [9], to food safety [10], agricultural productivity [11], and security [12].
Despite these notable achievements, e-nose technologies continue to face several key challenges that have slowed their broader adoption. These include: (i) sensor drift, which deteriorates performance over time [13]; (ii) cross-sensitivity, which leads to reduced specificity when sensors respond to multiple volatile compounds [14], in contrast to the modulation of the combinatorial code during the perception of a mixture of odorant observed in biology [15]; (iii) in the case of widely used metal oxide (MOx) sensing elements: significant power consumption due to high operating temperatures [16]; (iii) limited computational capacity for interpreting the complex combinatorial patterns inherent in real-world odors, as many e-nose systems rely on embedded microprocessors as the core control unit with limited processing capacity [17]; and (iv) insufficient adaptability, as current systems cannot dynamically refine their responses from experience, unlike their biological counterparts [6]. Observing how biological olfaction circumvents these problems suggests a promising model for the next generation of artificial olfactory devices, particularly given its robust signal processing capabilities, cross-selective receptor integration, and the ability to learn from repeated exposure [18, 19, 20].
@@
An important question is how olfactory receptor responses are distributed across different odorants. Until recently, consensus has been that individual receptors always respond to a range of volatile compounds [21, 22, 23, 24], implying that odors compete for overlapped receptor sites through a process often described as odor occlusion [25, 26]. Such competitive interactions can be highly non-linear: one odorant can amplify receptor activity, while another can dampen it, leading to outcomes that defy simple additive models [26]. However, as has recently been argued by Wachowiak et al. [27], these findings are mostly based on stimulation protocols which exceed naturally relevant concentration ranges by orders of magnitude. In fact, under naturally occurring concentrations, tuning curves of receptor neurons appear highly odorant-specific [28, 29, 30], in some cases activating only a single glomeruli [31]. Furthermore, a large meta analysis study [27] suggests that naturally occurring odor concentrations are often found near or below their detection threshold. Such view introduce many layers of complexity, making biological olfaction remarkably adept at navigating real-word odor spaces. Unlocking and understating those principle in engineering systems might push artificial olfaction beyond its current limits and toward robust, adaptable and context-aware sensing technologies.
@@ 1.1 Background & state-of-the-art
Metal oxide sensors, while effective for chemical detection, are limited both by their rather slow response time (on the order of many seconds), and drift over longer timescales [39, 40]. Drift manifests as temporal shifts in sensor response under identical working conditions, primarily caused by chemical and physical interactions at the sensing film microstructure, as well as external factors such as temperature and humidity variations. For example, recent analysis of the widely used gas sensor dataset by Vergara et al. [41] revealed how sensor drift can significantly affect recordings, hindering realistic assessment of sensor performance [39]. Although traditional approaches attempt to address drift through regression models and environmental compensation [42, 43], bio-inspired neuromorphic implementations offer promising alternatives through spike-based encoding [44] and adaptive learning mechanisms that could potentially provide stability and resistance to drift. For instance, Linster and Cleland in [45] model sensor inputs that are first normalized via inhibition in the glomerular layer, ensuring a concentration-invariant encoding. Synaptic plasticity based on spike-timing-dependent plasticity in the external plexiform layer then enables the system to continually adapt its internal representations of odors. This allows the network to refine its response profiles over time, compensating for both baseline shifts and changes in sensor responsiveness through experience-driven learning. Such neuromorphic approaches aim to reduce the complexity of the data and implement adaptive algorithms that can follow changing conditions, potentially offering a more robust solution to the persistent challenge of sensor drift.
@@
Building on these concepts, integrated analog on-chip learning circuitry has demonstrated efficacy in odor detection and classification. For example, Covington et al. in [35] used an 80 element chemoresistive microsensor array combined with a microfluidic package, while Tang et al. in [37] relied on commercially available sensors and a traditional signal processing pipeline in a microprocessor to distinguish three fruit fragrances. In a related work [36], Bernabei et al. utilized a large array of chemosensors with slightly varied sensor elements to emulate the overlap selectivity characteristic of biological olfaction, generating rich combinatorial responses. More recently, Rastogi et al. [46] introduced a neuromorphic analog e-nose that directly encodes the gas concentration using MOx sensors and an analog front-end that transforms sensor responses into spike-based representations. Their approach leverages the time difference between spikes in two distinct processing pathways to infer gas concentration, reducing data redundancy, and improving power efficiency.
@@
However, current polymer and metal oxide-based sensing elements remain relatively simple compared to biological olfactory systems, facing challenges with odor specificity and stability [47, 48]. Although these artificial sensors can detect certain volatile compounds, they often require high operating temperatures and exhibit limited effectiveness in discriminating complex molecular mixtures [16, 49]. A new generation of nanomaterials-based MOx sensors, and large-scale monolithically integrated nanotube sensor arrays [50], promises to alleviate some of these issues, but current implementations are still in their prototype phase [51, 14]. Moreover, their continuous sampling needs and high power consumption stand in stark contrast to the low-power, event-driven nature of biological olfaction [6]. These fundamental discrepancies in functionality and energy efficiency highlight the limitations of conventional chemosensors in achieving the richness and adaptability seen in biological odor detection.
@@
Notably, some more recent works have demonstrated a more complete reconstitution of the olfactory system, in particular, into bio-mimetic environments, for example membrane bound vesicles formed by membrane extraction, or virus-like particles [54, 55, 56, 57]. These approaches leverage the fact that insect olfaction is largely based on multivalent interactions between the conserved co-receptor Orco and the odorant binding subunits (odorant receptors, ORs), which together form a ligand-gated ion channel. Khadka et al. [56] have demonstrated successful insertion of multiple odorant receptors into surface-anchored lipid vesicles. Yamada et al. [57] have demonstrated parts per billion sensitivity of an Orco-OR complex inserted into a lipid bilayer to octenol, a biomarker in human breath. Detection of femtomolar concentrations of VUAA1, E2-hexenal, 4-ethylguaiacol molecules was demonstrated using odorant receptors Or10a, Or22a, and Or71a [56].
@@
The signal restoration challenge is being recognized as one of the central problems in neuromorphic olfaction [58, 59, 6]. In fact, for more than 20 years, researchers have concluded that the speed and accuracy of odor recognition in insects cannot be attributed solely to the performance of olfactory receptors due to their slow time constant and high variability. For instance, insects obtain fast, accurate odor recognition by transforming the slow, variable receptor signals in the antennal lobe by means of its network dynamics that converts these inputs into richer spatiotemporal patterns, especially during the initial transient, that carry far clearer, odor-specific information [60].
The requirement to correctly classify odor concentrations poses additional demand on the analyzing algorithms, as do signal occlusion and noise effects. Thus, several models of signal processing downstream of the sensor have been investigated. Examples of such models includes the architecture of the mammalian olfactory bulb [20], insect models that focus at the receptor level [61, 62], within the antennal lobe [60] or the macroglomerual complex [63]; and neuromorphic network that replicate insect coding [64, 65, 66] and learning [67] strategies. A comparative study even contrasts these biologically grounded approaches with classical signal processing strategies [68], highlighting the advantages of spatio-temporal codes for robust odor classification.
Given this rich literature we can roughly group methods for odor signal analysis into two classes: data analysis tools (e.g., principal component analysis, support vector machines) and biorealistic network models. While data analysis tools have been used for odor classification since many years [68, 61], biologically realistic network models are now studied with the goal of reaching the accuracy and efficiency of the animal olfactory system [63, 66, 20, 64]. Many of these models perform classification using a spatiotemporal code while learning in an unsupervised fashion. Additionally, some studies attribute the efficiency of natural systems to two main intrinsic mechanisms of biorealistic neural networks (i.e., Spiking Neural Network (SNN)): i) inhibition (which triggers a winner-takes-all competition [65, 63, 20]) and ii) temporal integration (due to neuron and network dynamics) [60, 20]. Recent work has also pursued the implementation of learning from teaching signals (supervised and reinforcement learning) in models of the insect olfactory system, in particular for associative learning [67, 69].
@@
Several hardware implementations have been demonstrated in digital Field Programmable Gate Arrays (FPGA) hardware [63], digital neuromorphic devices [20, 70] and analog neuromorphic chips [34, 65, 71, 46]. Recently, a bio-realistic network [20], based on the architecture of the mammalian olfactory bulb, was implemented on the Loihi neuromorphic digital chip, demonstrating the benefits of employing an algorithm based on online learning and spike timing. However, the system was trained using a separately recorded data set of pure chemicals [41], therefore lacking integration with real chemosensors, and criticism has been raised regarding the generalizability of the network model [39]. Beyond odor classification, recent developments have also tackled the challenge of odor localization; Schoepe et al. [70] demonstrated a fully event-driven neuromorphic implementation using stereo arrays of MOx sensors interfaced with a spiking neural network on SpiNNaker hardware [72]. Their approach effectively leveraged a spike-time-difference encoding strategy to infer odor source direction, while Dennler et al. [30] emphasized the effectiveness of neuromorphic event-driven sensing principles for efficiently capturing sparse and dynamic odor signals encountered in turbulent real-world environments.
The neuromorphic olfactory sensor literature indicates that there is considerable scope for improving neuromorphic chemosensory systems. There are several limitations in many of the current implementations, such as the lack of signal conditioning, the lack of integration of sensing and decoding in a single platform, high power consumption, and the poor sensitivity of the artificial sensor array. So far, the only truly integrated neuromorphic sensing platform for olfaction has been developed by Koickal et al. in 2006 [73, 34]. To our knowledge, this platform did not leave the proof-of-concept stage, and subsequent work by the authors has used conventional processing hardware [63]. Our present work is focused on integrating hardware realizations of all stages into one single system.
@@ 2 Co-design of synthetic biology, bio-inspired models, and neuromorphic systems
A comprehensive approach is crucial in developing advanced neuromorphic chemosensory systems that effectively integrate sensor arrays, decoding algorithms, and their physical realization [74]. In this work, we present a novel methodology and preliminary results on how to move beyond the State-Of-The-Art (SOTA) by addressing the lack of codesign between these three components. Unlike previous chemical detection systems inspired by the olfactory pathway, which have treated sensing [75, 61], signal processing [76, 77], and neuromorphic computation [20] as separate components, our approach integrates these aspects to tackle the challenge of chemosensory perception in a comprehensive way. In this work, we introduce a new concept to couple the depolarization of synthetic cells made from biological materials directly with CMOS electronic circuits designed for spike generation, we call these new class of bio-electronic devices hybrid synthetic sensory neurons, and we depict them in Fig. 1.
@@
In doing so, our framework harnesses the advantages of co-design in a joint environment. Here, the three disciplines (neuroscience, synthetic biology, neuromorphic engineering) are developed together and connected to each other (see Fig. 1):
@@
Our approach is not without challenges, as individual parts need to be optimized and tuned together but must still allow for the necessary flexibility. This flexibility is needed as knowledge and research about the individual parts progresses simultaneously. The parallel approach is, on one hand, due to constraints of scientific funding—collaborative grants typically work on a common timeline from proposal to project end. However, parallelized development is also, in general, desired for time efficiency. As such, the collaborative tools we develop, particularly those that bridge data-driven synthetic biology with electrical and information engineering, not only represent standalone contributions to each field, but might also be useful in other applications. In the following, we detail our co-design approach, validate our sensing pipeline using simulation-based modelling, and discuss ongoing challenges.
3 Results
@@ 3.1 Insect odor receptors provide a modular sensing platform
Insect olfaction is largely based on transmembrane protein complexes consisting of an odorant receptor (OR) and a co-receptor, Orco. While ORs diverge in their sequence and odor specificity between the multitude of insects and the ecological niches they inhabit, the co-receptor Orco is highly conserved across insect species. Several experiments show that it is possible to create functional receptor complexes by combining Orco of one species with a foreign OR [78, 79]. From an engineering point of view, this suggests the possibility of a modular platform in which the base assembly of an Orco-based sensing element allows the swapping of ORs with different response profiles to target various compounds. To demonstrate this modularity, we set out to computationally assemble Orco-OR complexes using AlphaFold modeling [80, 81].
@@
Figure 2: AlphaFold predictions suggest hundreds of possible stable complexes between D. melanogaster Orco and foreign odor receptors (OR). a1: AlphaFold ipTM score for Orco-OR complexes formed from D. melanogaster Orco combined with OR of different insect species (black), and combined with D. melanogaster (Dm)  OR (positive control, red) and M. musculus OR (negative control, grey). a2: Interface solvation energy prediction for Dm with foreign OR above the ipTM cut-off of 0.6. Dashed grey lines indicates cut-off thresholds. b1: Dm Orco (red) dimerization with M. sexta OR4 resulted in an ipTM score above 0.6 (dark grey) while the dimerization with L. migratoria OR113 led to a score below 0.6 (grey). b2:  Dm Orco and B. mori OR15 complex with favorable solvation energy (>5>-5> - 5 kcal/mol), whereas the dimerization with C. quinquefasciatus OR151 resulted in a low (<5<-5< - 5 kcal/mol) solvation energy. Varying number of bonds can be observed in the anchor domain (brackets). Green dashed lines indicate electrostatic interactions, while dashed blue lines represents hydrogen bonds.
@@
Specifically, we selected the Orco variant from the D. melanogaster (Dm) fruit fly and considered a total of 348 Orco-OR complexes, including non-Dm ORs. Foreign ORs derived from either other insect species or from a mammal (mouse, M. musculus) as a negative control. Each predicted Orco-OR structure was evaluated using the interface-predicted template modeling (ipTM) score, with higher values indicating greater confidence in the prediction accuracy of the protein-protein interface. An empirical cut-off point of 0.6 was chosen (Figure 2a1). The protein-protein interface refers to the region where the two proteins directly interact with each other. The interface of the Alphafold-predicted Orco-OR complexes is of great interest as it provides insight into the plausibility and structural integrity of the possible receptor complex. [82, 81]. As expected, the positive control, consisting only of Orco-OR complexes from the same species (Dm), yielded mostly acceptable modeling outputs, as reflected in ipTM scores above 0.6. In contrast, the negative control, which used the mammalian OR that are not expected to form complexes with Orco, did not produce any acceptable models (Figure 2a1). Interestingly, for Dm Orco and ORs from foreign species, we found a large fraction of potentially accurate models with ipTM scores above the 0.6 cut-off. This suggests that AlphaFold models are relevant for understanding the structures of Dm Orco in complex with ORs from foreign species.
@@
To further evaluate the structures, we considered a second metric: interface solvation energy, which provides insight into the stability of the non-covalent binding between Orco and OR (Figure 2a2). More negative values indicate a more favorable binding interface [83]. We found that the obtained values for foreign OR were similar to those of the native Dm Orco-OR complexes (Fig. 2b2), indicating comparable complex stability even for foreign ORs. Furthermore, we manually inspected the most favorable predictions for each group. In both cases, the interactions between the two subunits were confined to a segment of helical bundles stabilized by hydrogen bonds and electrostatic interactions located away the membrane interface (Fig. 2a2), consistent with the anchor domain and interactions identified in experimental studies [84, 85, 78]. These results suggest that the OR structures generated by AlphaFold are promising candidates for computational exploration of the wide diversity of ORs.
@@
Figure 3: Comparison of ligand binding energies of Dm - Or69a, Or98a, Or19a and Or92a. Each datapoint corresponds to one odor molecule. Left: Terpene subset. Right: All odors from the DoOR dataset with neuron response above the spontaneous firing rate (SFR). Dashed line shows linear regression with $R^{2}=0.2$.
To demonstrate the potential to build a digital representation of OR-odor interactions, we challenged the generated AlphaFold Orco-OR structures with in-vivo recordings of neuronal activity. We used the DoOR 2.0.1 dataset [86, 24], a comprehensive database of olfactory responses from various measurement techniques such as single sensillum recordings, calcium imaging, and heterologous expression [87, 88, 22, 89, 21, 90, 86, 91, 92, 93, 94, 95, 96, 97, 98, 99, 62, 100, 101, 102, 103]. In the DoOR dataset, all recordings have been merged and normalized across studies, providing response data for 78 types of sensory units to 692 odorants. Both excitatory and inhibitory responses are included in the dataset, represented as responses above or below the spontaneous firing rate (SFR). The magnitude of the SFR varies between ORs and is typically below 0.1 normalized units. As we aim to reconstitute OR, we exclude units from the dataset which can not be clearly linked to a single OR type, such as responses from multiple receptor types or ionotropic receptors. This leaves a total of 52 possible OR types as the starting point of our analysis. Note that the DoOR dataset contains recordings from stimulations at high concentrations not found in nature. As discussed in Section 1 and Ref. [27], odor occlusion and separability are strongly affected by odor concentration, which will need to be taken into account in real-world deployment of our sensing platform (see Discussion & Outlook).
@@
For each OR-odor combination, the DoOR dataset provides a normalized in-vivo neuron response (‘consensus dataset with global normalization’ in [24]). If the predicted OR structures are accurate, we expected that neuronal activity would correlate with the binding of the response-inducing molecule to the OR model. We initially focused on terpenes, a class of carbohydrates with a large natural abundance, because we speculated that these molecules were a feasible target for a computationally relatively simple molecular docking approach. We calculated the binding energy of terpenes, to the generated OR structures using Vina Dock (see Materials and Methods) for a subset of Dm ORs and plotted these against the measured neuronal response from the DoOR dataset (Fig. 3, left). We found that the difference in binding energy between responding (> 0) and non-responding neurons ($\leq$ 0) was highly significant (Student’s t-test, $p<3\cdot 10^{-7}$), suggesting that the predicted structures have realistic binding pockets for terpenes. Motivated by this, we investigated the same Dm OR subset using all molecules present in the DoOR dataset. Here, we obtained a weak correlation between neuron activation and binding energy (Fig. 3, right). At this point it is unclear whether the scatter can be explained by the uncertainty of the molecular docking method or point to more complex molecular recognition mechanisms. Currently, we are investigating an even larger number of ORs and binding molecules with refined computational approaches. If the responses of olfactory neurons to odors are indeed explained simply by the binding energy of an odor to the OR, the molecular recognition process of insect olfaction can be condensed into a measure of this binding energy. Taken together, these results show the potential to use computational tools to predict OR to odor response.